Biot savart law for a straight wire

Grade:12

1 Answers

Neeti
8 years ago
Bio Savart law gives us the value of magnetic field for a straight conductor at a point P at a distance r from the wire. 
it is given as 
B = k x[idlsin\theta /r2]
where K is \mu / 4\pi = 10-7